nitric oxide?

What mass of CaH3N can be made when 21.6 g of CaH6 react with 21.6 g of


1
Expert's answer
2022-05-05T12:26:03-0400

4C3H6(g) + 6NO(g) → 4C3H3N(s) + 6H2O(l) + N2(g)


n(C3H6)=m(C3H6)/M(C3H6) =21.6 g/42.08 g/mol=0.513 mol

n(NO)=m(NO)/M(NO) =21.6 g/30.01 g/mol = 0.720 mol

n(C3H6)/4=0.128 mol;

n(NO)/6=0.120 mol

So NO is the limiting reactant.

n(C3H3N)=4/6 n(NO)=0.480 mol

m(C3H3N)=n(C3H3N)×M(C3H3N)=0.480 mol ×53.06 g/mol = 25.5 g